A beam LO weighing 50N rests against a wall, the lower part of the beam extending 15 m from the base of the wall. Calculate the minimum value of the force F that can be applied to keep the beam in equilibrium. The coefficient of friction at L and O is 0.30.



1
Expert's answer
2021-01-27T09:31:28-0500

Answer

Consider that beam is making angle 45° with horizontal so

Moment about point O

"50\times7.5\times \sqrt{2}\times sin45+0.3N'\times\sqrt{2}cos45=N'\times15"


Therefore

N'=35.7Newton.


Force that requires to keep the beam in

equilibrium

"35.7\times0.3+Q=50cos45^°"


So

Q=24.6N